# Coherent-state boundary conditions as the first-principles origin of background fields in QED

### v2

@@ Abstract
QED formulated in prescribed classical background electromagnetic fields is a standard framework for strong-field and laser–matter interactions. It is usually treated as a theory modified by externally imposed fields, obscuring its precise relation to full QED and, in particular, the role of asymptotic boundary conditions for the gauge field.
Starting from an operator-based formulation, we show that QED with background fields is not a distinct theory but arises as a well-defined boundary-condition limit of full QED, in which the classical background field emerges from coherent-state boundary conditions on the quantized electromagnetic field. In this limit, the conventional generating functional used in calculations with prescribed background fields is recovered naturally, while relaxing the boundary conditions allows depletion and backreaction effects to be incorporated within the same framework. The central new result is a rigorous operator-level proof of the equivalence between the fixed background-field approximation and coherent-state asymptotic boundary conditions—a formulation that, to our knowledge, has not been made explicit in previous approaches.
We further demonstrate that the apparent time dependence of background fields does not originate from an explicitly time-dependent Hamiltonian, but instead reflects the choice of picture—Schrödinger versus Heisenberg—in the underlying quantum theory. Rather than introducing new properties of coherent states, our analysis provides a first-principles reinterpretation of the fixed background-field approximation as a controlled and picture-dependent limit of full QED.
322-4 Oroshi-cho, Toki, Gifu, 509-5292, Japan
February 17, 2026
@@ 1 Introduction
Classical background electromagnetic (EM) fields are a central ingredient in quantum electrodynamics (QED), underpinning a wide range of applications from laser–matter interactions to strong-field and external-field problems [32, 21, 29, 14, 31, 4, 6]. In these approaches, the background field is treated as a prescribed classical quantity, while quantum fluctuations of the EM and matter fields are quantized on top of it. In this sense, one works within a formulation of QED in which classical and quantized EM fields coexist. This framework has proven extraordinarily successful in practice, enabling non-perturbative treatments of intense fields and the description of phenomena such as nonlinear Compton scattering [31, 3, 27, 28], Breit–Wheeler pair production [31, 27, 28, 30], and vacuum polarization effects [14, 7].
@@
Despite this success, the conceptual status of the fixed background field approximation remains subtle. In this paper, we use the term “background field approximation” to denote formulations of QED in which the EM field is decomposed into a prescribed classical background and quantized fluctuations, with the background treated as fixed and unaffected by the quantum dynamics. In conventional formulations—often summarized under the heading of QED in external EM fields or the Furry picture [9, 25]—the background field is introduced as an explicit c-number input, frequently endowed with spacetime dependence from the outset. While operationally well defined, this procedure leaves open a fundamental question: how does a classical background EM field arise from the underlying quantum theory of the field, and in what precise sense is the fixed background an approximation?
@@
A natural bridge between fully quantized gauge fields and classical backgrounds is provided by coherent states of the radiation field [13, 12, 37]. Coherent states possess classical-like expectation values while remaining genuine quantum states, and have long been recognized as an appropriate description of idealized laser fields [34, 23]. It is also well known that scattering amplitudes involving coherent states can be reexpressed as amplitudes computed in the presence of background fields, a correspondence discussed in various contexts since early work such as Ref. [8] and in later analyses of backreaction and depletion effects [19]. These observations suggest that the role of background fields in scattering processes may ultimately be tied to the choice of asymptotic boundary conditions for the quantized EM field.
@@
The central claim of this paper is that QED with prescribed classical background fields is not a distinct or modified theory. Rather, the fixed background field approximation can be identified unambiguously as a controlled boundary-condition limit of full QED. We show that QED with prescribed background fields arises naturally as a limiting case of scattering between coherent states of the quantized EM field, corresponding to fixed coherent-state boundary conditions. In this sense, the background field is not a fundamental input, but an emergent classical object whose status and limitations can be precisely identified.
@@
Our formulation is developed at the operator level, starting from the Schrödinger and Heisenberg pictures of QED [20, 33]. This perspective makes it possible to disentangle two conceptually distinct issues that are often conflated in standard treatments. First, the classical background field is identified with the expectation value of the quantized gauge field in a coherent state, subject to appropriate physical-state conditions, here implemented via the Gupta–Bleuler formalism [15, 2] or, equivalently, BRST cohomology [1, 38, 22]. Second, the spacetime dependence of the background field—commonly taken as an external input—is shown to arise from the choice of picture: while the Hamiltonian remains time independent in the Schrödinger picture, the Heisenberg-picture evolution of operators induces a time-dependent background expectation value. In particular, the time dependence of background fields is not introduced ad hoc at the level of the Hamiltonian, but emerges naturally from the underlying quantum dynamics.
@@
Within this framework, the conventional generating functional of QED with a prescribed background field [14, 36] is recovered as a special case corresponding to fixed coherent-state boundary conditions. Allowing for transitions between different coherent states provides a systematic and fully quantum-mechanical description of effects beyond the fixed background field approximation, such as depletion and backreaction [19], without double counting of photon degrees of freedom. For clarity, we refer to this operator-based formulation—in which classical background fields arise as coherent-state expectation values subject to fixed coherent-state asymptotic boundary conditions—as coherent-field QED. More recently, coherent states have been employed as a technical ingredient to relate amplitudes on background fields to gravitational amplitudes via the classical double copy [18]. Our aim here is different: rather than exploiting this equivalence as a computational tool, we reinterpret the fixed background field approximation itself as a limiting procedure based on coherent-state boundary conditions. While operator-based approaches to QED with prescribed background fields—most notably the unstable-vacuum formalism of Fradkin, Gitman and Shvartsman [7], and related analyses by Gavrilov and Gitman [11]—demonstrate that coherent states can mimic prescribed classical fields in specially chosen initial states, these works formulate the background as an externally imposed c-number field and treat the coherent-state displacement operator within the interaction picture. Consequently, the origin of the background field is picture-dependent and not tied to a boundary-condition prescription in full QED. In contrast, the present work identifies the formulation of QED with prescribed background fields itself as a well-defined coherent-state asymptotic boundary-condition limit, formulated at the operator level and independent of picture choice. This allows a first-principles reinterpretation of QED with prescribed background fields and clarifies its precise relation to the underlying quantum theory.
@@ 2 Gupta–Bleuler Condition and Coherent Laser EM Fields
We restrict attention to coherent states satisfying the Gupta–Bleuler condition, thereby defining physically admissible coherent laser states. In the following discussion the EM field $\hat{A}_{\mathrm{H}}=\hat{A}_{(+)\mathrm{H}}+\hat{A}_{(-)\mathrm{H}}$ (The subscript “H” refers objects in the Heisenberg picture) is assumed to satisfy the Gupta–Bleuler condition [15, 2]. Hatted symbols denote operators. Suppose that $\hat{A}_{(+)\mathrm{H}}$ and $\hat{A}_{(-)\mathrm{H}}$ are the positive-frequency and negative-frequency components of $\hat{A}_{\mathrm{H}}$. The Gupta–Bleuler condition is imposed as
@@
and that conjugate relation is $\langle\mathrm{Physical}_{\mathrm{H}}|\partial_{\mu}\hat{A}_{(-)\mathrm{H}}^{\mu}(x)=0$, where $|\mathrm{Physical}_{\mathrm{H}}\rangle$ denotes any physical state. In this work we adapt the Gupta–Bleuler condition to a coherent state in the Heisenberg picture, thereby defining a coherent laser EM field. A coherent state [13, 12, 37, 34, 23] is specified by the following eigen equation
@@
where, $\hat{\mathbb{I}}$ is the identity operator [12, 23, 8]. In the subsequent analysis, this section provides the operator-based definition of a coherent laser EM field.
3 Basic formalism
@@
Related operator-based approaches to QED with prescribed background fields using coherent-state boundary conditions were discussed previously, notably in Ref. [7, 11]. In those formulations, asymptotic coherent states are introduced via a displacement operator described as time “independent,” implicitly within an interaction- or asymptotic-picture setting. Here, we make the role of picture dependence explicit by showing how the time dependence of the background field arises from operator evolution in the Heisenberg picture while remaining absent in the Schrödinger picture by construction.
In strong-field QED/nonlinear QED experiments at high-intensity laser facilities [10, 24, 35], scattering processes are typically formulated starting from a Hamiltonian of the form $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$], where the laser field $\mathcal{A}$ is introduced as a prescribed classical background [14, 31, 4, 25, 16]. It is then commonly stated that the theory can be reformulated in the Furry picture, in which the background field becomes explicitly time dependent and is incorporated into the free Hamiltonian. From the viewpoint of the Schrödinger picture, however, the background field $\mathcal{A}$ is a classical c-number quantity without intrinsic time dependence, while in the Furry picture it appears as a time-dependent external field [31, 4, 6, 9, 25]. In conventional treatments, the origin of this time dependence is not derived but assumed, which obscures how the background field is related to the underlying quantum dynamics. The detailed construction of the Furry picture in coherent-field QED will be developed in Sect. 4.
@@ 3.1 Departure from the Schrödinger picture
We begin by recalling that, in the Schrödinger picture, field operators are time independent, while the time evolution is entirely carried by the quantum states [20, 33]. Within this picture, the introduction of a coherent state via a displacement operator provides a natural and unambiguous definition of a classical background field as the expectation value of a quantized EM field. Importantly, this construction is performed at the operator level and does not require modifying the Hamiltonian or introducing an explicitly time-dependent external field by hand.
We consider QED in the presence of a coherent background EM field and reformulate it in terms of coherent-field QED. To this end, we assume that a coherent background laser field is always present. Particles other than the coherent EM field are denoted abstractly by $|\Psi_{\mathrm{S}}(t)\rangle$, where, the subscript S denotes the Schrödinger picture. Thus, in the Schrödinger picture, the state containing both the coherent laser field and other particles can be written as
@@
Here $t_{0}$ is chosen as the time at which the Schrödinger and Heisenberg pictures are matched. Although one would usually write $\hat{O}_{\mathrm{S}}(\boldsymbol{x})=\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})$, here we retain the explicit notation. Equation (16) is easily obtained from Eq. (7) via the common relation of an observable $\hat{O}_{\mathrm{S}}=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{D}_{\mathrm{S}}(\alpha)$, that is, $\hat{O}_{\mathrm{S}}=\hat{O}_{\mathrm{H}}(t_{0})$. Importantly, the background field $\mathcal{A}$ automatically satisfies the Lorenz gauge condition (5) via the Gupta–Bleuler condition (1) [15, 2] evaluated at the reference time $t_{0}$. Since we are in the Schrödinger picture [20, 33], the Hamiltonian $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] is time-independent. We therefore write
@@
The key point of this subsection is that the usual formulation of QED in prescribed background fields emerges directly from the coherent-field QED framework once the coherent state is factored out in the Schrödinger picture. In particular, starting from the fully quantum description $|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ and $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]$, the displacement by a coherent state produces the familiar fixed-background description $|\Psi_{\mathrm{S}}(t)\rangle$ and $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] widely used in strong-field QED. This shows that the fixed-background field approximation is not the starting point but rather a derived limit within coherent-field QED. Moreover, because depletion and path-integral formulations require transitions between different coherent states, the fundamental structure of this work is the underlying coherent-field QED structure defined by
@@
not its fixed-background reduction.

### v3

# 322-4 Oroshi-cho, Toki, Gifu, 509-5292, Japan

@@ Abstract
QED formulated in prescribed classical background electromagnetic fields is a standard framework for strong-field and laser–matter interactions. It is usually treated as a theory modified by externally imposed fields, obscuring its precise relation to full QED and, in particular, the role of asymptotic boundary conditions for the gauge field. Starting from an operator-based formulation, we show that QED with background fields is not a distinct theory but arises as a well-defined boundary-condition limit of full QED, in which the classical background field emerges from coherent-state boundary conditions on the quantized electromagnetic field. In this limit, the conventional generating functional used in calculations with prescribed background fields is recovered naturally, while relaxing the boundary conditions allows depletion and backreaction effects to be incorporated within the same framework. The central new result is a rigorous operator-level proof of the equivalence between the fixed background-field approximation and coherent-state asymptotic boundary conditions—a formulation that, to our knowledge, has not been made explicit in previous approaches. We further demonstrate that the apparent time dependence of background fields does not originate from an explicitly time-dependent Hamiltonian, but instead reflects the choice of picture—Schrödinger versus Heisenberg— in the underlying quantum theory. Rather than introducing new properties of coherent states, our analysis provides a first-principles reinterpretation of the fixed background-field approximation as a controlled and picture-dependent limit of full QED.
@@ 1 Introduction
Classical background electromagnetic (EM) fields are a central ingredient in quantum electrodynamics (QED), underpinning a wide range of applications from laser–matter interactions to strong-field and external-field problems Saiamin et al. (2006); Joachain et al. (2011); Raicher et al. (2014); Greiner et al. (1985); Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023). In these approaches, the background field is treated as a prescribed classical quantity, while quantum fluctuations of the EM and matter fields are quantized on top of it. In this sense, one works within a formulation of QED in which classical and quantized EM fields coexist. This framework has proven extraordinarily successful in practice, enabling nonperturbative treatments of intense fields and the description of phenomena such as nonlinear Compton scattering Ritus (1985); Brown and Kibble (1964); Nikishov and Ritus (1964, 1967), Breit–Wheeler pair production Ritus (1985); Nikishov and Ritus (1964, 1967); Reiss (1962), and vacuum polarization effects Greiner et al. (1985); Fradkin et al. (1991).
@@
Despite this success, the conceptual status of the fixed background-field approximation remains subtle. In this paper, we use the term “fixed background-field approximation” to denote formulations of QED in which the EM field is decomposed into a prescribed classical background and quantized fluctuations, with the background treated as fixed and unaffected by the quantum dynamics. In conventional formulations—often summarized under the heading of QED in external EM fields or the Furry picture Furry (1951); Moortgat-Pick (2009)—the background field is introduced as an explicit c-number input, frequently endowed with spacetime dependence from the outset. While operationally well defined, this procedure leaves open a fundamental question: how does a classical background EM field arise from the underlying quantum theory of the field, and in what precise sense is the fixed background an approximation?
@@
A natural bridge between fully quantized gauge fields and classical backgrounds is provided by coherent states of the radiation field Glauber (1963b, a); Sudarshan (1963). Coherent states possess classical-like expectation values while remaining genuine quantum states, and have long been recognized as an appropriate description of idealized laser fields Sargent III et al. (1974); Loudon (2000). It is also well known that scattering amplitudes involving coherent states can be reexpressed as amplitudes computed in the presence of background fields, a correspondence discussed in various contexts since early work such as Ref. Frantz (1965) and in later analyses of backreaction and depletion effects Ilderton and Seipt (2018). These observations suggest that the role of background fields in scattering processes may ultimately be tied to the choice of asymptotic boundary conditions for the quantized EM field.
@@
The central claim of this paper is that QED with prescribed classical background fields is not a distinct or modified theory. Rather, the fixed background-field approximation can be identified unambiguously as a controlled boundary-condition limit of full QED. We show that QED with prescribed background fields arises naturally as a limiting case of scattering between coherent states of the quantized EM field, corresponding to fixed coherent-state boundary conditions. In this sense, the background field is not a fundamental input, but an emergent classical object whose status and limitations can be precisely identified.
@@
Our formulation is developed at the operator level, starting from the Schrödinger and Heisenberg pictures of QED Itzykson and Zuber (1980); Sakurai (1967). This perspective makes it possible to disentangle two conceptually distinct issues that are often conflated in standard treatments. First, the classical background field is identified with the expectation value of the quantized gauge field in a coherent state, subject to appropriate physical-state conditions, here implemented via the Gupta–Bleuler formalism Gupta (1950); Bleuler (1950) or, equivalently, BRST cohomology Becchi et al. (1976); Tyutin (1975); Kugo and Ojima (1979). Second, the spacetime dependence of the background field—commonly taken as an external input—is shown to arise from the choice of picture: while the Hamiltonian remains time independent in the Schrödinger picture, the Heisenberg-picture evolution of operators induces a time-dependent background expectation value. In particular, the time dependence of background fields is not introduced ad hoc at the level of the Hamiltonian, but emerges naturally from the underlying quantum dynamics.
@@
Within this framework, the conventional generating functional of QED with a prescribed background field Greiner et al. (1985); Schwinger (1951) is recovered as a special case corresponding to fixed coherent-state boundary conditions. Allowing for transitions between different coherent states provides a systematic and fully quantum-mechanical description of effects beyond the fixed background-field approximation, such as depletion and backreaction Ilderton and Seipt (2018), without double counting of photon degrees of freedom. For clarity, we refer to this operator-based formulation—in which classical background fields arise as coherent-state expectation values subject to fixed coherent-state asymptotic boundary conditions—as coherent-field QED. More recently, coherent states have been employed as a technical ingredient to relate amplitudes on background fields to gravitational amplitudes via the classical double copy Ilderton and Lindved (2025). Our aim here is different: rather than exploiting this equivalence as a computational tool, we reinterpret the fixed background-field approximation itself as a limiting procedure based on coherent-state boundary conditions. While operator-based approaches to QED with prescribed background fields—most notably the unstable-vacuum formalism of Fradkin, Gitman and Shvartsman Fradkin et al. (1991), and related analyses by Gavrilov and Gitman Gavrilov and Gitman (1990)—demonstrate that coherent states can mimic prescribed classical fields in specially chosen initial states, these works formulate the background as an externally imposed c-number field and treat the coherent-state displacement operator within the interaction picture. Consequently, the origin of the background field is picture-dependent and not tied to a boundary-condition prescription in full QED. In contrast, the present work identifies the formulation of QED with prescribed background fields itself as a well-defined coherent-state asymptotic boundary-condition limit, formulated at the operator level and independent of picture choice. This allows a first-principles reinterpretation of QED with prescribed background fields and clarifies its precise relation to the underlying quantum theory.
@@ 2 Gupta–Bleuler Condition and Coherent Laser EM Fields
We restrict attention to coherent states satisfying the Gupta–Bleuler condition, thereby defining physically admissible coherent laser states. In the following discussion the EM field $\hat{A}_{\mathrm{H}}=\hat{A}_{(+)\mathrm{H}}+\hat{A}_{(-)\mathrm{H}}$ (The subscript “H” refers to objects in the Heisenberg picture) is assumed to satisfy the Gupta–Bleuler condition Gupta (1950); Bleuler (1950). Hatted symbols denote operators. Suppose that $\hat{A}_{(+)\mathrm{H}}$ and $\hat{A}_{(-)\mathrm{H}}$ are the positive-frequency and negative-frequency components of $\hat{A}_{\mathrm{H}}$. The Gupta–Bleuler condition is imposed as
@@
and that conjugate relation is $\langle\mathrm{Physical}_{\mathrm{H}}|\partial_{\mu}\hat{A}_{(-)\mathrm{H}}^{\mu}(x)=0$, where $|\mathrm{Physical}_{\mathrm{H}}\rangle$ denotes any physical state. In this work we adopt the Gupta–Bleuler condition for a coherent state in the Heisenberg picture, thereby defining a coherent laser EM field. A coherent state Glauber (1963b, a); Sudarshan (1963); Sargent III et al. (1974); Loudon (2000) is specified by the following eigen equation
@@
where, $\hat{\mathbb{I}}$ is the identity operator Glauber (1963a); Loudon (2000); Frantz (1965). In the subsequent analysis, this section provides the operator-based definition of a coherent laser EM field.
3 Basic formalism of coherent-field QED
@@
Related operator-based approaches to QED with prescribed background fields using coherent-state boundary conditions were discussed previously, notably in Ref. Fradkin et al. (1991); Gavrilov and Gitman (1990). In those formulations, asymptotic coherent states are introduced via a displacement operator described as time “independent,” implicitly within an interaction- or asymptotic-picture setting. Here, we make the role of picture dependence explicit by showing how the time dependence of the background field arises from operator evolution in the Heisenberg picture while remaining absent in the Schrödinger picture by construction.
In strong-field QED/nonlinear QED experiments at high-intensity laser facilities Gales et al. (2018); Mirzaie et al. (2024); Sarri et al. (2025), scattering processes are typically formulated starting from a Hamiltonian of the form $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$], where the laser field $\mathcal{A}$ is introduced as a prescribed classical background Greiner et al. (1985); Ritus (1985); Di Piazza et al. (2012); Moortgat-Pick (2009); Hartin (2018). It is then commonly stated that the theory can be reformulated in the Furry picture, in which the background field becomes explicitly time dependent and is incorporated into the free Hamiltonian. From the viewpoint of the Schrödinger picture, however, the background field $\mathcal{A}$ is a classical c-number quantity without intrinsic time dependence, while in the Furry picture it appears as a time-dependent external field Ritus (1985); Di Piazza et al. (2012); Fedotov et al. (2023); Furry (1951); Moortgat-Pick (2009). In conventional treatments, the origin of this time dependence is not derived but assumed, which obscures how the background field is related to the underlying quantum dynamics. The detailed construction of the Furry picture in coherent-field QED will be developed in Sect. 4.
@@ 3.1 Departure from the Schrödinger picture
We begin by recalling that, in the Schrödinger picture, field operators are time independent, while the time evolution is entirely carried by the quantum states Itzykson and Zuber (1980); Sakurai (1967). Within this picture, the introduction of a coherent state via a displacement operator provides a natural and unambiguous definition of a classical background field as the expectation value of a quantized EM field. Importantly, this construction is performed at the operator level and does not require modifying the Hamiltonian or introducing an explicitly time-dependent external field by hand.
We consider QED in the presence of a coherent background EM field and reformulate it in terms of coherent-field QED. To this end, we assume that a coherent background laser field is included from the outset. Particles other than the coherent EM field are denoted abstractly by $|\Psi_{\mathrm{S}}(t)\rangle$, where, the subscript S denotes the Schrödinger picture. Thus, in the Schrödinger picture, the state containing both the coherent laser field and other particles can be written as
@@
Here $t_{0}$ is chosen as the time at which the Schrödinger and Heisenberg pictures are matched. Although one would usually write $\hat{O}_{\mathrm{S}}(\boldsymbol{x})=\hat{O}_{\mathrm{S}}(t_{0},\boldsymbol{x})$, here we retain the explicit notation. Equation (16) is easily obtained from Eq. (7) via the common relation of an observable $\hat{O}_{\mathrm{S}}=\hat{D}_{\mathrm{S}}^{\dagger}(\alpha)\hat{A}_{\mathrm{S}}(t_{0},\boldsymbol{x})\hat{D}_{\mathrm{S}}(\alpha)$, that is, $\hat{O}_{\mathrm{S}}=\hat{O}_{\mathrm{H}}(t_{0})$. Importantly, the background field $\mathcal{A}$ automatically satisfies the Lorenz gauge condition (5) via the Gupta–Bleuler condition (1) Gupta (1950); Bleuler (1950) evaluated at the reference time $t_{0}$. Since we are in the Schrödinger picture Itzykson and Zuber (1980); Sakurai (1967), the Hamiltonian $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] is time-independent. We therefore write
@@
The key point of this subsection is that the usual formulation of QED in prescribed background fields emerges directly from the coherent-field QED framework once the coherent state is factored out in the Schrödinger picture. In particular, starting from the fully quantum description $|\alpha_{\mathrm{S}}(t),\Psi_{\mathrm{S}}(t)\rangle$ and $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}]$, the displacement by a coherent state produces the familiar fixed background description $|\Psi_{\mathrm{S}}(t)\rangle$ and $H_{\mathrm{QED}}[\hat{\psi}_{\mathrm{S}},\hat{\bar{\psi}}_{\mathrm{S}},\hat{A}_{\mathrm{S}}+\mathcal{A}\hat{\mathbb{I}}$] widely used in strong-field QED. This shows that the fixed background-field approximation is not the starting point but rather a derived limit within coherent-field QED. Moreover, because depletion and path-integral formulations require transitions between different coherent states, the fundamental structure of this work is the underlying coherent-field QED structure defined by
@@
not its fixed background reduction.